Dess:

A supermarket finds that the number of boxes of a new soda drink sold increases each week. In the first week, only 24 boxes of the soda drink were sold. In the next week, 56 boxes of the soda drink were sold and in the third week 88 boxes of the soda drink were sold. The number of boxes of soda drink sold each week represents an arithmetic sequence. What is the explicit rule for the arithmetic sequence that defines the number of boxes of soda drink sold in week n?

Allison:

an = 24 + 32(n-1)

Allison:

56 boxes, d = a2 - a1 = 56-24 = 32

Allison:

an = a1 + d*(n-1) where a1 is the starting value, d is the common difference, and n is the nth term in the sequence. In this problem, an represents the number of boxes sold in week n. The starting value is a1 = 24 boxes. The common difference in obtained by subtracting any two successive terms. Since a1 = 24 boxes and a2 = 56 boxes, d = a2 - a1 = 56-24 = 32

Dess:

Thank you.

Allison:

Anytime. Please close this question when you're done. :)

Warriorz13:

Esh sexeh shorts TinyAlly

Falconmaster:

omg stop